An impure sample of Pb(NO3)2 weighing 2.420 g was completely dissolved
in water. When all of the Pb+2 was precipitated from this solution with H2SO4 it produced 0.626 g of PbSO4(s). What was the percentage of Pb(NO3)2 in the sample?
1
Expert's answer
2020-09-14T08:25:46-0400

The equation for the reaction is

"Pb(NO_3)_2+H_2SO_4\to PbSO_4+2HNO_3"

The ratio of "Pb(NO_3)_2"

"1" atom of lead to "2" atoms of nitrogen to "6" atoms of Oxygen.

Simplified to "1Pb^{2+} to 2NO_3^{-}"

The molecular weight of "Pb(NO_3)_2=331.2g\/mol"

Percentage yield"=" "Actual yield\over theoretical yield" "\times 100"

Maximum yield of "Pb(NO_3)_2" "=" "2.420g\times" "207.2g\over (331.2\times 2)g" "=0.756g"

Percentage yield "=" "0.626g\over 0.756g" "\times" "100"

"=82.80\%"